OpenStudy (anonymous):

Find an equation for the quadratic function described: The function f has zeros 5 and 1 and f(0)=1. THANKS SO MUCH

OpenStudy (anonymous):

function with zeros at 5 and 1 is f(x) = a ( x - 5 ) ( x - 1 ) = f(x) = ax^2 - 6ax + 5a = f (0) = 5a = 1 = a = 1/5 Hence , f(x) = 1/5 x^2 - 6/5 x + 1

OpenStudy (anonymous):

Thank you!